# Anchored Learning for On-the-Fly Adaptation — Extended Technical Report

### v2

@@
Abstract
This study presents “anchor critics”, a novel strategy for enhancing the robustness of reinforcement learning (RL) agents in crossing the sim-to-real gap. While RL agents can be successfully trained in simulation, they often encounter difficulties such as unpredictability, inefficient power consumption, and operational failures when deployed in real-world scenarios. We identify that naïve fine-tuning approaches lead to catastrophic forgetting, where policies maintain high rewards on frequently encountered states but lose performance on rarer, yet critical scenarios. Our method maximizes multiple Q-values across domains, ensuring high performance in both simulation and reality. Evaluations demonstrate that our approach enables behavior retention in sim-to-sim gymnasium tasks and in sim-to-real scenarios with racing quadrotors, achieving a near-50% reduction in power consumption while maintaining controllable, stable flight. We also contribute SwaNNFlight, an open-source firmware for testing adaptation techniques on real robots. Video: https://youtu.be/WgDFaS5nXrw
@@ I Introduction
Reinforcement Learning (RL) has emerged as a powerful approach for solving complex control problems where traditional techniques may be inadequate. Due to the expensive and potentially risky nature of real-world robot interaction, RL agents are typically trained and evaluated in simulated environments. However, recent work has shown that controllers applied to real systems can exhibit undesirable properties, from minor control issues to severe hardware failures [1, 2, 3]. These problems stem from the challenge of crossing the domain gap between simulated training environments and the real world—the “reality gap”.
While direct on-robot training raises safety and cost concerns, making simulated training attractive, the transition from simulation to reality presents significant challenges[4, 5]. Real-world control scenarios can generate distributions of trajectories where critical scenarios occur only in the tails. This makes the expected discounted sum of rewards, and thus, the associated Q-value, an inadequate measure of agent performance across important tasks. Recent works have identified that traditional fine-tuning approaches often result in catastrophic forgetting, where policies lose their previously well-behaved performance when adapting to new environments [6, 7]. Indeed, we observed this phenomenon across every benchmark task we tested.
To address these challenges, we propose a method that maximizes multiple Q-values across domains in a multi-objective manner. By leveraging the flexibility of simulated training, we can control the data distribution by evaluating agents on critical scenarios, allowing the Q-value from simulation to serve as effective “anchors” for key performance metrics while adapting to Q-values learned from real-world data. This approach enables robust sim-to-real transfer while preserving essential behaviors learned in simulation.
Our work also addresses the practical challenges of training RL agents online using resource-constrained robotic platforms. Many existing solutions offload computations to ground-station servers, limiting the robots’ autonomy and practicality due to network dependencies and communication constraints [8, 9]. We present SwaNNFlight, an open-source firmware enabling on-board inference with efficient ground station communication for policy updates, making our approach viable for real-world applications requiring live adaptation and real-time control responses.
@@ II Related Work
SwaNNFlight extends Neuroflight, an open-source firmware for NN-based flight control [10, 11]. While Neuroflight demonstrated the potential of embedding NN models into flight control systems, it faced challenges such as high training variance, poor sim-to-real transfer, and limited flexibility in controller updates. We address these issues by tackling the problem of adaptating RL agents between domains and by providing a flexible firmware for live controller updates.
A body of work exists on facing adaptation challenges:
  • $\bullet$
    Model-based methods aim to identify optimal parameters from observations [12, 13, 14], but struggle with accurately modeling complex real-world dynamics.
  • $\bullet$
    Model-free approaches like Meta RL [15, 16, 17] train policies to adapt to different environments, but are limited by the fidelity of simulated environments in representing real-world dynamics shifts.
  • $\bullet$
    Trust Region Policy Optimization (TRPO) [18] and Proximal Policy Optimization (PPO) [19] improve stability by constraining policy changes during optimization. Our anchor critics method shares the goal of preserving existing behavior but focuses on optimizing performance across both simulated and real environments.
  • $\bullet$
    Liquid Neural Networks [20, 21, 22, 23, 24] build adaptability into the network structure itself, showing promise in control tasks with smaller models. However, their effectiveness with out-of-distribution data needs further exploration.
  • $\bullet$
    Some approaches employ RL directly on robotic platforms [25, 26], addressing safety concerns through constrained exploration, though this relies on safety criteria modeling.
  • $\bullet$
    Hebbian plasticity approaches offer adaptive control in soft robotics, addressing morphological damage and improving human-robot interactions. Recent work has applied these techniques to compliant robotics, meta-learning for damage adaptation, and improving handshaking synchronization [27, 28, 29].
It’s important to note that systematic comparison of these approaches is challenging due to the lack of standardized benchmarks for evaluating domain adaptability in reinforcement learning. This underscores the need for more unified evaluation methods in this rapidly evolving field.
@@ III Live Adaptation with Anchor Critics
The main problem our work seeks to address is that of transferring learned control policies between similar domains in real-time. In our main use-case, this is transferring controllers trained in simulation to the real-world, where simulated dynamics may not be sufficiently accurately captured to enable sim-to-real transfer without performance degradation. A common issue of sim-to-real policy transfer is erratic control, which leads to high power consumption and mechanical stress. Our proposed solution attempts to bridge the domain gap during transfer by adapting controllers to the new domain during live operation, aiming to refine and enhance the control response.
@@
The aforementioned ‘domain gap’ is characterized by a change in the underlying dynamics between two domains. Attempting to simply adapt policies learned in simulation to new data derived from the new dynamics was observed to quickly lead to catastrophic forgetting where ‘good’ behavior learned in simulation would be forgotten, causing the policies to become unstable and unpredictable. [6, 7] When new observations are not consistent with simulated observations, unpredictably aberrant behavior is a safety risk. Unlike in simulation, it is not possible in real-life to safely initialize the control system in arbitrary and diverse conditions. Experience in the new domain needs to start being gathered from a ‘safe’ region and boundaries of safety can be expanded as the policy adapts to the new domain. However, this induces a bias in the adapted policy towards those initial ‘safe’ experiences – and while some degree of bias is beneficial for the policy to adapt itself, balancing the effects of the new experience against the beneficial experiences gained in simulation can be tricky.
@@ A Case Against Mixing Experience Buffers During Adaptation
It may seem reasonable, as a first-pass solution, to adapt policies by mixing simulated and real experience during fine-tuning. However, this can be problematic in two critical ways. (i) It breaks the Markov assumption by introducing unobserved hidden variables relating to which dynamical environment any given sample of experience may have come from, thus introducing uncertainty and inconsistency to the relationship between states, actions, and rewards – where consistency is critical for effective RL training. (ii) It would be difficult to ensure a balance in the diversity of the two disparate datasets. This could introduce a reward skew, as discussed further in Section III-A. Instead, we suggest that separating the value functions that drive policy optimization between domains would enable weighing how much policies ’care’ about one domain over the other. Thus allowing a tunable trade-off of stability over adaptability and vice versa.
@@ III-A Data Distribution Impact on Policy Optimization
Multi-objective optimization is significantly impacted by the coverage of the data upon which solutions are optimized. Consider, for example, a two-objective policy optimization which aims to simultaneously minimize actuation (analogous to power consumption), $\phi_{p}$, and goal-tracking error, $\phi_{e}$, such that the optimal policy, $\pi^{*}$ is given by:
@@ III-B Illustrating Reward Skew
This is the maximum possible value, as $R_{d}$ is always near 1 in this scenario. The final Q-value function becomes
$Q^{\pi}(s,a)\approx w_{p}Q_{p}^{\pi}(s,a)+\frac{w_{d}}{1-\gamma}.$ (7)
In this case, the optimal policy $\pi^{*}$ will be:
$\pi^{*}=\operatorname*{arg\,max}_{\pi}w_{p}V_{p}^{\pi}(s).$ (8)
@@
This optimal policy will balance both power efficiency and trajectory tracking, depending on the relative weights $w_{p}$ and $w_{d}$.
@@ III-C Anchor Critics
Our results suggest that the issue of catastrophic forgetting naturally arises in a wide range of robot controllers when live adaptation with RL is naïvely performed. We observed catastrophic forgetting within only a few updates across various benchmark control problems, underscoring the severity of the issue. In simulation, it is possible to constantly train for a wide-range of maneuvers without risking physical safety. Conversely, many aspects of real robot data may be uncontrollable. Valid live adaptation must incorporate real data while respecting the behavior learned in simulation. We term the domain in which the agent is trained the source domain and the domain in which the agent is adapted the target domain. In order to tackle the issues presented, we propose adapting agents on two main value criteria: (i) the Q-value $Q_{\pi}$ on observed on data from the target and (ii) the Q-value $Q_{\Psi}$ learned on data from the source. We term the latter the ‘anchor’ critic as it serves to anchor the policy behavior to the behavior learned in simulation.
We then treat these Q-values as smooth constraints that must be jointly satisfied. By requiring that step-rewards are normalized—i.e., $R:s\times a\to[0,1]$—we can bound the maximum possible Q-value as follows:
$Q_{\max}=\mathbf{E}_{\tau}[R(\tau)=1]=\sum^{\infty}_{k=0}\gamma^{k}=\frac{1}{1-\gamma}.$ (10)
By dividing the Q-values by $Q_{\max}$ we obtain a normalized Q-value. Treated as a constraint, these normalized Q-values capture how much the current policy satisfies our goals. Since we would like to satisfy multiple Q-values, we then aim to maximize what is known as the product T-norm in Fuzzy Logic. We motivate this choice in Section VII.
@@
The anchored policy optimization goal, given anchor critic $Q_{\Psi}$, extends the classical actor-critic policy loss $J_{\pi_{\theta}}^{\Psi}$ as
$-\mathbb{E}_{s_{r}\sim\mathcal{D}_{\text{T}},s_{s}\sim\mathcal{D}_{\text{S}}}[Q_{\pi_{\theta}}(s_{r},\pi_{\theta}(s_{r}))\cdot Q_{\Psi}(s_{s},\pi_{\theta}(s_{s}))^{w_{\Psi}}]$ (11)
where $\mathcal{D}_{\text{T}}$ and $\mathcal{D}_{\text{S}}$ are the replay buffers for the target and source domains respectively, and $w_{\Psi}$ is the anchor weighting. The amount of weight offered to the anchor dictates the trade-off between satisfying performance objectives in source and target domains.
@@
To keep anchors representative on the source domain, we keep updating $Q_{\Psi}$ during training, following the typical actor-critic update. For this, we use the current adapted policy but only on data sampled from the replay buffer $\mathcal{D}_{\text{S}}$ generated during source domain training. The Q-value $Q_{\pi_{\theta}}$ is conversely trained only on data sampled from the target domain replay buffer $\mathcal{D}_{\text{T}}$.
@@ IV Anchor Critics in Simulation
Following this, we conduct sim-to-sim transfer evaluations using modified Gymnasium environments in Section IV-C. These modifications are reflect specific adaptation goals, allowing us to showcase the prevalence of catastrophic forgetting in traditional approaches and demonstrate how anchor critics effectively mitigate this issue. We provide anchors as a library and allow the experiments to be reproduced [30].
@@ IV-A Testing Anchors on Inverted Pendulum
To test the general viability of anchor critics, we implemented them in three popular contemporary RL algorithms: DDPG [31], SAC [32], and TD3 [33] in Fig. 1 and tested them on a modified inverted pendulum problem. The typical goal of the task is to stabilize a torque-controlled pendulum vertically upright, but we instead aim to hold the pendulum at a specified angle. As shown in Fig. 2, naively fine-tuned, the agents learn exclusively solve the task in the source or target domain. On the other hand, anchors enable agents to learn to compromise as shown in Fig. 2(c). The evolution of the pendulum angle is shown in Fig. 3. Additional data for more agents and for agents trained with different algorithms are available in Section VIII.

### v3

# Sim-Anchored Learning for On-the-Fly Adaptation

@@
Sim-Anchored Learning for On-the-Fly Adaptation
Bassel El Mabsout1, Siddharth Mysore1, Shahin Roozkhosh1, Kate Saenko1 and Renato Mancuso1 1Affiliated with the Computer Science department at Boston University
@@ I Introduction
Reinforcement Learning (RL) offers powerful solutions for complex control problems where traditional techniques may be inadequate. Due to the expensive and potentially risky nature of real-world robot interaction, RL agents are typically trained and evaluated in simulated environments. However, recent work has shown that controllers applied to real systems can exhibit undesirable properties, from minor control issues to severe hardware failures [1, 2, 3]. These problems stem from the challenge of crossing the domain gap between simulated training environments and the real world—the “reality gap”.
While direct on-robot training raises safety and cost concerns, making simulated training attractive, the transition from simulation to reality presents significant challenges[4, 5]. While the ”reality gap” often focuses on dynamics discrepancies, a key related difficulty arises from the distributional sim-to-real gap. Real-world experience collection, especially during initial adaptation, frequently yields limited or skewed state distributions compared to the broader coverage achievable in simulation. Indeed, real-world control scenarios often generate distributions where critical scenarios occur only in the tails. Relying solely on such skewed real-world data means the learned RL value function (e.g., the Q-value) often fails to capture the designer’s intended priorities. While reward functions are typically considered the primary specification of intent, we posit that comprehensive priorities are also implicitly encoded within the simulation through controllable choices like task setups and state initializations. This broader, simulation-encoded intent is precisely what risks being lost during adaptation on limited real-world data, leading to catastrophic forgetting. We propose using the value function derived from this intentionally designed simulation as an “anchor for intentionality” during real-world adaptation. Optimizing against both the anchor critic ($Q_{\Psi}$) and a real-world critic ($Q_{\pi}$) allows adaptation while preserving crucial, intentionally designed behaviors. This facilitates robust sim-to-real transfer that respects the intended operational goals across the full state space.
Implementing and evaluating adaptation strategies like anchor critics directly on resource-constrained robots presents significant practical challenges. While many existing solutions offload computations to ground-station servers [6, 7], this can limit autonomy and real-time responsiveness. To facilitate the real-world validation of our anchor critic approach and enable practical on-the-fly adaptation, we developed and contribute SwaNNFlight: an open-source firmware stack. By enabling efficient on-board inference coupled with ground-station communication for policy updates, SwaNNFlight provides the necessary infrastructure for the live adaptation experiments discussed herein, making robust, real-time control adjustments viable for real-world robotic applications.
@@ II Related Work
Bridging the sim-to-real gap in RL requires addressing both dynamics discrepancies and the critical distributional shifts between simulation and reality highlighted previously [4, 5]. Existing transfer strategies approach this challenge in various ways, but often struggle to adapt effectively without losing the comprehensive behavioral intent encoded within the simulation’s design.
Sim-to-Real Transfer Strategies: Techniques like domain randomization aim to improve robustness by varying simulation parameters, though fully capturing real-world variability can be challenging [4]. Domain adaptation methods fine-tune policies on real data but often suffer from catastrophic forgetting, where agents overfit to the limited target distribution and lose crucial behaviors learned across the simulation’s broader context [8]. Our work specifically targets this forgetting problem arising from distributional shifts during adaptation.
Stable Adaptation Techniques: Ensuring stability during adaptation is crucial. Methods like Trust Region Policy Optimization (TRPO) and Proximal Policy Optimization (PPO) [9] limit policy updates to prevent divergence. However, these methods do not explicitly focus on preserving the specific, intentionally designed behaviors learned from the diverse scenarios in simulation. Our Anchor Critics approach complements stability concerns by directly anchoring the policy to the simulation-encoded value function ($Q_{\Psi}$) while adapting based on real-world experience ($Q_{\pi}$).
On-Robot Adaptation and Infrastructure: Applying these strategies on resource-constrained robots presents practical hurdles. Prior work like Neuroflight demonstrated embedding NN controllers in firmware but faced challenges with sim-to-real transfer and update flexibility [10, 11]. Our SwaNNFlight firmware stack builds upon this, enabling efficient on-board inference coupled with ground-station communication for live policy updates [12]. This provides the necessary platform for deploying advanced adaptation methods like Anchor Critics in real-time.
Positioning Relative to Prior Work: While existing work addresses dynamics gaps and stability, our method uniquely tackles catastrophic forgetting induced by distributional shifts by anchoring adaptation to the simulation’s comprehensive design intent. Combined with the SwaNNFlight infrastructure, we offer a practical and theoretically grounded solution for robust live adaptation.
@@ III Live Adaptation with Anchor Critics
The main problem our work seeks to address is that of transferring learned control policies between similar domains in real-time. In our main use-case, this is transferring controllers trained in simulation to the real-world, where simulated dynamics may not be sufficiently accurately captured to enable sim-to-real transfer without performance degradation. A common issue of sim-to-real policy transfer is erratic control, which leads to high power consumption and motor over-stress. Our proposed solution attempts to bridge the domain gap during transfer by adapting controllers to the new domain during live operation, aiming to refine and enhance the control response.
@@
The ‘domain gap’ involves dynamics changes, and attempting adaptation often leads to catastrophic forgetting where simulation-learned behavior is lost, causing instability [8, 13]. Furthermore, the distribution of states encountered during real-world operation can be highly skewed, potentially underrepresenting critical scenarios (like an emergency maneuver). Optimizing solely on this skewed distribution risks learning policies that forget the broader priorities encoded across the simulation’s design (including tasks, initializations, and rewards). When new observations are not consistent with simulated observations, unpredictably aberrant behavior is a safety risk. Unlike in simulation, it is not possible in real-life to safely initialize the control system in arbitrary and diverse conditions. Experience in the new domain needs to start being gathered from a ‘safe’ region and boundaries of safety can be expanded as the policy adapts to the new domain. However, this induces a bias in the adapted policy towards those initial ‘safe’ experiences – and while some degree of bias is beneficial for the policy to adapt itself, balancing the effects of the new experience against the beneficial experiences gained in simulation can be tricky.
@@ A Case Against Mixing Experience Buffers During Adaptation
It may seem reasonable, as a first-pass solution, to adapt policies by mixing simulated and real experience during fine-tuning. However, this can be problematic in two critical ways. (i) It breaks the Markov assumption via unobserved hidden variables about the data’s domain origin, introducing inconsistency. (ii) Ensuring a meaningful balance between potentially skewed real-world data and intentionally curated simulation data is difficult, risking reward skew (Section III-A). Instead, we suggest that separating the value functions that drive policy optimization between domains would enable weighing how much policies ’care’ about one domain over the other. Thus allowing a tunable trade-off of stability over adaptability and vice versa.
@@ III-A Data Distribution Impact on Policy Optimization
Multi-objective optimization is significantly impacted by the coverage of the data upon which solutions are optimized. The value of simulation lies partly in the designer’s ability to enforce broad data coverage, reflecting intended priorities through controllable choices like task setups and initializations, a luxury often unavailable in real-world data collection. Consider, for example, a two-objective policy optimization which aims to simultaneously minimize actuation (analogous to power consumption), $\phi_{p}$, and goal-tracking error, $\phi_{e}$, such that the optimal policy, $\pi^{*}$ is given by:
@@ III-B Illustrating Reward Skew
This is the maximum possible value, as $R_{d}$ is always near 1 in this scenario. The final Q-value and optimal policy becomes
$Q^{\pi}(s,a)$ $\approx w_{p}Q_{p}^{\pi}(s,a)+\frac{w_{d}}{1-\gamma}$ (7)
$\pi^{*}$ $=\operatorname*{arg\,max}_{\pi}w_{p}V_{p}^{\pi}(s).$
@@
This optimal policy will balance both power efficiency and trajectory tracking, depending on the relative weights $w_{p}$ and $w_{d}$. This example highlights how reliance on skewed data can misdirect policy optimization. Our proposed anchor critic mechanism, detailed next, directly addresses this by incorporating the value function derived from an intentionally balanced simulation design, thereby preserving sensitivity to all intended objectives.
@@ III-C Anchor Critics
Our results suggest that the issue of catastrophic forgetting naturally arises in a wide range of robot controllers when live adaptation with RL is naïvely performed. We observed catastrophic forgetting within only a few updates across various benchmark control problems, underscoring the severity of the issue. Simulation provides a controlled environment where diverse maneuvers can be safely trained, allowing for the design of tasks, rewards, and state distributions that collectively encode the desired operational intent. Conversely, real-world data collection may be limited by safety constraints and yield skewed distributions unrepresentative of the full intended operating range. Valid live adaptation must therefore incorporate real data while respecting the comprehensive behavioral profile captured by the simulation’s design. We term the domain in which the agent is initially trained (typically simulation) the source domain and the domain in which the agent is adapted (typically reality) the target domain. To tackle the challenges presented, we propose adapting agents based on two main value criteria: (i) the Q-value $Q_{\pi}$ learned on data observed from the target domain, reflecting adaptation to current dynamics and the observed data distribution, and (ii) the Q-value $Q_{\Psi}$ learned on data from the source domain, representing the performance according to the simulation-encoded intent. We term the critic providing $Q_{\Psi}$ the ‘anchor’ critic, as it anchors the policy optimization to the intended behavior profile shaped by the deliberate simulation design choices (tasks, initializations, rewards), preventing deviations caused by potentially limited or skewed target domain data.
We formulate the adaptation challenge as a multi-objective problem: the agent must simultaneously satisfy performance objectives derived from the source domain’s design intent and objectives learned from the target domain’s experiences. Following Fulfillment Priority Logic (FPL) [14], we represent these objectives as fulfillments in the range $[0,1]$. Assuming rewards are normalized such that Q-values can be appropriately scaled to represent expected fulfillment within $[0,1]$ (e.g., $Q(s,a)\times(1-\gamma)$), we can apply FPL composition operators. For notational simplicity, we will use $Q$ to denote these fulfillment-compatible values representing the expected performance in the source ($Q_{\Psi}$) and target ($Q_{\pi_{\theta}}$) domains. Our goal is to achieve high simultaneous fulfillment across both domains. To compose these Q-values for joint satisfaction, we utilize the FPL conjunction operator $\boldsymbol{\land}^{0}$ (geometric mean), as defined in FPL [14], applied to the target fulfillment $Q_{\pi_{\theta}}$ and the anchor fulfillment raised to the priority weight $w_{\Psi}$, i.e., $Q_{\Psi}^{w_{\Psi}}$. The geometric mean, a specific case ($p=0$) of FPL’s power mean conjunction, inherently rewards joint satisfaction—the value is high only if both $Q_{\pi_{\theta}}$ and the prioritized $Q_{\Psi}$ are high—and implicitly prioritizes the improvement of less-fulfilled objectives, consistent with FPL’s goal of balancing competing goals.
@@
The anchored policy optimization goal, given anchor critic $Q_{\Psi}$, thus modifies the standard actor loss $J_{\pi_{\theta}}$ to maximize this geometric mean composition $J_{\pi_{\theta}}^{\Psi}$:
$\mathbb{E}_{s_{r}\sim\mathcal{D}_{\text{T}},s_{s}\sim\mathcal{D}_{\text{S}}}[Q_{\pi_{\theta}}(s_{r},\pi_{\theta}(s_{r}))\boldsymbol{\land}^{0}(Q_{\Psi}(s_{s},\pi_{\theta}(s_{s}))^{w_{\Psi}})]$ (9)
where $\mathcal{D}_{\text{T}}$ and $\mathcal{D}_{\text{S}}$ are the replay buffers for the target and source domains respectively. The exponent $w_{\Psi}$ acts as a user-defined priority weight. While FPL uses a specific syntax ($[\phi]_{\delta}$) for explicit prioritization, $w_{\Psi}$ serves the analogous purpose here, allowing the user to control the relative importance of fulfilling the source domain’s design intent ($Q_{\Psi}$) versus the target domain’s observed objectives ($Q_{\pi_{\theta}}$). A higher $w_{\Psi}$ prioritizes preserving the original simulation behavior.
@@
To keep anchors representative on the source domain, we keep updating $Q_{\Psi}$ (and thus $Q_{\Psi}$) during training, following the typical actor-critic update but using only data sampled from $\mathcal{D}_{\text{S}}$. The Q-value $Q_{\pi_{\theta}}$ (and $Q_{\pi_{\theta}}$) is conversely trained only on data sampled from $\mathcal{D}_{\text{T}}$.
@@ IV Anchor Critics in Simulation
Following this, we conduct sim-to-sim transfer evaluations using modified Gymnasium environments in Section IV-C. These modifications, which introduce shifts in dynamics or task distributions, reflect specific adaptation goals. They allow us to showcase how such distributional shifts can induce catastrophic forgetting even without a reality gap, and demonstrate how anchor critics effectively mitigate this issue by preserving intended behaviors learned in the source domain. We provide anchors as a library and allow the experiments to be reproduced [12].
@@ IV-A Testing Anchors on Inverted Pendulum
To test the general viability of anchor critics, we implemented them in three popular contemporary RL algorithms: DDPG [15], SAC [16], and TD3 [17] in Fig. 1 and tested them on a modified inverted pendulum problem. The typical goal of the task is to stabilize a torque-controlled pendulum vertically upright, but we instead aim to hold the pendulum at a specified angle. As shown in Fig. 2, the naively fine-tuned agent exclusively solves the target domain task, forgetting the source behavior. The anchor critic, however, represents the value function associated with the source domain’s intended behavior (leaning left). By incorporating this anchor, the agent is encouraged to find a policy that balances the conflicting target objective (leaning right) with the original intent, resulting in a compromise behavior (leaning straight up) as shown in Fig. 2(c). The evolution of the pendulum angle under anchored adaptation is shown in Fig. 3. Additional data for more agents and for agents trained with different algorithms are available in Section VIII.